Write the slope intercept form equation of the line passing through (2, -3) and (-2, -5)

Answer by stanbon(75887)

Write the slope intercept form equation of the line passing through (2, -3) and (-2, -5)
------
y + 3 = [(-3--5)/(2--2)](x-2)
------
y+3 = (2/4)(x-2)
y = (1/2)x -1 -3
y = (1/2)x-4
--------------
Cheers,
Stan H.